OpenStudy (anonymous):

PLEASE HELP!!! An expression is shown below: square root of 32 plus square root of 2 Which statement is true about the expression? It is rational and equal to 4. It is rational and equal to 5. It is irrational and equal to 5 multiplied by square root of 2. It is irrational and equal to 4 multiplied by square root of 2.

OpenStudy (anonymous):

@Hero @satellite73

OpenStudy (anonymous):

\[\sqrt{32}+\sqrt2\]?

OpenStudy (anonymous):

since \(32=16\times 2\) you have \[\sqrt{32}=\sqrt{16\times 2}=\sqrt{16}\sqrt2=4\sqrt2\]

OpenStudy (anonymous):

then \[4\sqrt2+\sqrt2=5\sqrt2\]

OpenStudy (anonymous):

as usual, the answer is C it is always C

OpenStudy (anonymous):

THANK YOU SO MUCH!!!!! :)

OpenStudy (anonymous):

yw